OpenStudy (anonymous):

What is the equation of a line passing through the point (a, b) and having a slope of b?

OpenStudy (anonymous):

A. y = −bx − b(1 − a) B. y = bx − b(a − 1) C. \[y=\frac{ 1 }{ b}x+b(a-1)\] D. \[y=-\frac{ 1 }{ b }x+b(1-a)\]

OpenStudy (nirmalnema):

(y-y1) = m (x-x1) (y-b)=b(x-a) y = b(x-a) + b y = bx - ba + b y = bx - b(a - 1)